Show by an example that in general a continuous function is neither convex nor concave


1
Expert's answer
2021-03-31T13:54:38-0400

Take the sine function. Then take any two points a,b from "[0,\pi]." Then in this range the function is convex. The reason is "f(x)=sin \ x\Rightarrow f^{''}(x)=-sin \ x <0." Since in the given range the sine function is positive. Again from "[\pi,2\pi ], f^{''}(x)>0" since in the given range the sin function is negative.Hence in that range the function is concave. Hence in general a continuous function is neither concave or convex.